OpenStudy (anonymous):

The office jet printer can copy janet's dissertation in 16 minutes. The laser jet can copy the same document in 14 minutes. If the two machines work together, how long would it take to copy the dissertation?

OpenStudy (anonymous):

Start with this: (1/16)+(1/14)=(1/x)

OpenStudy (anonymous):

I cannot remeber if i have to add or multiply

OpenStudy (anonymous):

you add, just like i have written it out

OpenStudy (anonymous):

2/30

OpenStudy (anonymous):

1/30 or 30 minutes???

OpenStudy (anonymous):

i got : 224/30 or 7.5 minutes

OpenStudy (anonymous):

are you sure you evluated the equation i gave you?

OpenStudy (anonymous):

use product /sum: (14*16)/(14+18)=224/30=7.5